# Tree Attention: Topology-aware Decoding for Long-Context Attention on GPU clusters

### v1

@@ Abstract
Self-attention is the core mathematical operation of modern transformer architectures and is also a significant computational bottleneck due to its quadratic complexity in the sequence length. In this work, we derive the scalar energy function whose gradient computes the self-attention block, thus elucidating the theoretical underpinnings of self-attention, providing a Bayesian interpretation of the operation and linking it closely with energy-based models such as Hopfield Networks. Moreover, due to this formulation, we discover that we can use efficient and optimized automatic-differentiation techniques to derive a highly efficient Tree Attention algorithm to compute the gradient of the energy and hence self-attention. Our formulation reveals that the reduction across the sequence axis can be efficiently computed in parallel through a tree reduction. Our algorithm, for parallelizing attention computation across multiple GPUs, enables cross-device decoding to be performed asymptotically faster (up to $8\times$ faster) than alternative approaches such as Ring Attention, while also requiring significantly less communication volume and incurring $2\times$ less peak memory. Our code is publicly available here: https://github.com/Zyphra/tree_attention.
@@ 1.  Introduction
During inference, the attention block largely determine the computational and memory requirements, which become more demanding as the input sequence length increases. Although LLMs generate one token at a time, the entire sequence of past tokens must still be stored in memory and used to compute attention scores during generation. Since attention performs a similarity matching of every token representation with every other, it incurs quadratic computational complexity in terms of flops. Moreover, it requires the entire sequence to be stored in memory during autoregressive generation.
@@
While there has been recent advances in extending LLMs to extremely long contexts (up to 1M tokens)Chen et al. (2023); kai (2023); Peng et al. (2023), inferencing at such contexts lengths at scale still remains extremely expensive. This expense is unfortunate since long context models allow for qualitatively new capabilities such as extremely large scale in-context learning entire small datasets held in the prompt Reid et al. (2024); Lee et al. (2024); Bertsch et al. (2024), natural direct access to multimodal continuous data which is currently intractable for current transformers and which therefore must currently undergo a lossy tokenization scheme Reid et al. (2024); Team (2024), and potentially enables removing tokenization altogether and operating directly at the byte level Xue et al. (2022); Wu et al. (2024).
@@
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu and Dao (2023); Dao and Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024). Alternately, other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory efficient attention Rabe and Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works, Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. This is the regime of primary interest of this paper. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
@@
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices instead of linearly as alternatives such as Ring Attention Liu et al. (2023). Our topology aware approach illustrated in Fig. 1. significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
@@
2.  Related Works
Following the ubiquitous success of the transformer architecture, there has been significant effort to try to mathematically understand the nature and meaning of the attention operation and link it to existing and related literatures such as recurrent networks and Hopfield Networks. A number of recent works have attempted to study self-attention mathematically through the lens of energy functions. Ramsauer et al. (2020) pioneered this field by performing a similar but distinct analysis to relate self-attention with the modern Hopfield networks Krotov and Hopfield (2016), thus providing a novel and insightful interpretation of self-attention as performing hetero-associative memory lookups using a high-powered nonlinear similarity function Krotov (2021); Millidge et al. (2022). This work was later extended by Hoover et al. (2023) who derived a modified version of the transformer based off an energy function. Beyond this, other recent works attempt a Bayesian reformulation of attention by deriving a probabilistic generative model which matches the operations performed in a self-attention operation Singh and Buckley (2023).
However, while, it has long been known that the softmax operation can be derived as the gradient of the following scalar function:
@@ 2.  Related Works
known as the log-sum-exp, deriving an equivalent function for the self-attention block has not yet been performed.
In particular, consider the energy function proposed in Hoover et al. (2023) which was the logsumexp. Since the gradient in the update rule of that paper is taken with respect to the input to the block, the resulting function is a modified version of the self attention operation. Similarly, the update rule in Ramsauer et al. (2020) requires the tying of certain weights (K and V) in the attention operation. This restricts the Hopfield derivation to modelling autoassociative retrieval networks, while transformer attention is heteroassociative.
To advance beyond this, our energy function depends on a source tensor $\zeta$ (see section 4.), an auxiliary vector, and the gradient we take is with respect to this tensor. We will observe that this leads to the exact self attention operation when we take the gradient. Consequently, our formulation enables us to express self attention with no weight tying.
@@ 3.  Self-Attention
The self attention operation can be represented as a set of dot product similarity searches between queries and keys. These similarity scores are then reduced along the sequence axis and softmaxed, so that for a given query, there is a probability distribution of the similarities of each given key. We then take the expectation of the value vectors against this distribution. We denote the queries assigned to a sequence of length $N$ as $\{q_{a},a=1,\cdots N\}$, where each query is a vector of size $d$ that stands for hidden dimension: $q_{a}\in\mathbb{R}^{d}$ and similarly the keys and values $\{(k_{a},v_{a}),a=1,\cdots N\}$. Attention can be written as,
@@
The dot product, more explicitly reads:
@@
The capital Latin indices ($A$) span the hidden dimension $d_{h}$. Naively computing attention in this way requires materializing the $qk$ matrix with quadratic computational and memory cost in the sequence length.
Memory efficient attention Rabe and Staats (2021) proposes an iterative way to compute the softmax similarities without ever having to materialize the full attention matrix. It performs the following operations, one query (or a chunk of queries) at a time:
@@
Then, once the values $v$ and softmax denominator $d$ are computed we can divide to get the final softmaxed scores $z^{(j)}=\frac{n^{(j)}}{d^{(j)}}$ for every query index $j$. Since we compute attention in this iterative manner, the memory requirements are significantly reduced.
Flash attention Dao et al. (2022), utilizes a similar approach to reduce the memory and computational cost of attention but the algorithm is not adapted for multi-GPUs. Flash attention performs the iterative algorithm of Rabe and Staats (2021) in a blockwise manner, utilizing the block-parallel computational primitives available inside single GPU tensorcores. Additionally, it precisely sizes the blocks such that they can fit into the SRAM of the GPU for the entire attention computation, effectively performing kernel fusion and preventing many unnecessary IO operations.
4.  Self Attention as a gradient
Here, we will show how the self-attention operation can be written as the gradient of a certain scalar function. In particular, we will define a scalar function that depends on the keys, queries, values and additionally, on an auxiliary vector we will refer to as the source tensor $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self attention operation. We need the source in order to write down the generating function of the moments of the distribution above. The source is the variable with respect to which we can Taylor expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
@@ 4.  Self Attention as a gradient
This terminology is inspired by work on energy-based models in machine learning Beal (2003); LeCun et al. (2006); Song and Kingma (2021). A summary of variables and indices is provided in appendix  11.
@@ 4.1.  The energy function
In what follows, we will show how the energy function we are after is given by the generating function for the cumulants of the distribution given by attention scores. Taking inspiration from statistical mechanics where an analogous cumulant generating function defines the Helmholtz Free energy Wikipedia (2024), we dub our cumulant generating function the energy function for self attention.
Let us focus on the case where there is a single query only. As noted above, we leverage the fact that the attention operation can be seen as the computation of the expectation value of the vectors $v$ in the distribution set by the attention scores $z$:
@@
We see that the probability density is given by:
@@
Typically, the denominator or normalization factor is identified with the so called partition function:
@@
Now, we can compute any moment of the distribution by as the $n$th Taylor coefficient of $Z(\zeta)$ $\forall A_{1},A_{2},\cdots\in\{1,\cdots,d_{h}\}$ :
@@
So we see that the first moment can be written as:
@@
This quantity is what we call the generating function, or the free energy:
@@
To compute causal self attention, then we introduce $N$ sources $\zeta^{i}$ each $\in\mathbb{R}^{d}$ and take
@@
[…]
@@
The generalization to the multi-head attention case is straightforward. There we have one key, query and value per head. For $n_{h}$ total heads, the generating function takes the form:
@@
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We will proceed further just focusing on the single head case as it makes the presentation more simple and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix 10.
@@ 4.2.  Bayesian interpretation
The fact that it is possible to derive the self-attention operation as the minimization of an energy function implies that it is possible to provide a Bayesian gloss on self-attention by identifying a likelihood function and show that we can obtain the forward pass of the attention block from computing the maximum a posteriori estimate of this likelihood.
@@
We denote by $x$ the input to the self attention block from which we obtain $q,k,v$ from multiplying it by the weights $W_{Q},W_{K},W_{V}$ respectively. Let us minimize the above with respect to $\zeta$ and $z$ simultaneously:
@@ 5.  Tree Attention
In this section we will show how the formulation of the attention operation as the gradient of an energy function suggests an efficient parallel strategy for computing it. The key insight is to leverage an efficient algorithm to compute the energy, and then differentiate it in order to obtain an efficient algorithm to compute attention.
@@ 5.1.  Computing the energy efficiently
A crucial and important fact is that both $\textrm{logsumexp}_{a}$ and $\max_{a}$ are associative operations:
@@ 5.2.  Efficient parallel decoding
This algorithm requires three Allreduce operations in total which means that the time complexity it requires is $O(3(N/p+\log p))$.
@@ 5.3.  Implementation and topology-awareness
(a) Execution time of of 16 head Tree Attention vs Ring Attention for different sizes of GPU cluster (from 1 to 16 H100 DGX nodes). Relative Execution Times are indexed to the Ring Attention times at a sequence length of 80k tokens.
@@
In our empirical experiments , we use Flash Attention 2 Dao (2023) within each device, both for our algorithm and for Ring Attention [^1] A JAX based Ring Attention implementation that uses Flash Attention 2 can be found here: https://github.com/nshepperd/flash_attn_jax.. We provide a simple JAX implementation of our method in Appendix 8.. Note that our method mirrors Flash Decoding fla (2024) except in that case, the parallelization happens at the level of different streaming multiprocessors (SMs) within a GPU whereas we parallelize between different GPUs. All computations are performed in BF16.
@@ 6.  Results
Similar to Ring Attention, Tree Attention is an exact computation of Attention. Since training and evaluation metrics are the same as Attention, our experimental results are focused primarily on latency in section 6.1., peak memory usage in section 6.2. and communication volumes in section 6.3.. Since our algorithm computes numerically identical results as the forward pass of standard Attention, our performance results transfer seamlessly to Transformer architectures.
@@ 7.  Discussion and Conclusion
Theoretical asymptotic improvements should be demonstrable in the single-device case, but in practice, this advantage is not realized. SMs in H100 and older architectures communicate through shared global memory, not in a peer-to-peer fashion like individual GPUs in a multi-device setup, making the reduced communication negligible. However, recent experimental instructions in the H100 have enabled peer-to-peer SM communication  NVIDIA (2024), suggesting that clever use of these instructions could lead to speedups over Flash Attention on a single device.

### v2

@@ Abstract
Self-attention is the core mathematical operation of modern transformer architectures and is also a significant computational bottleneck due to its quadratic complexity in the sequence length. In this work, we derive the scalar energy function whose gradient computes the self-attention block, thus elucidating the theoretical underpinnings of self-attention, providing a Bayesian interpretation of the operation and linking it closely with energy-based models such as Hopfield Networks. Our formulation reveals that the reduction across the sequence axis can be efficiently computed in parallel through a tree reduction. Our algorithm, for parallelizing attention computation across multiple GPUs enables cross-device decoding to be performed asymptotically faster (up to $8\times$ faster in our experiments) than alternative approaches such as Ring Attention, while also requiring significantly less communication volume and incurring $2\times$ less peak memory. Our code is publicly available here: https://github.com/Zyphra/tree_attention.
@@ 1.  Introduction
During inference, the attention block largely determines the computational and memory requirements, which become more demanding as the input sequence length increases. Although LLMs generate one token at a time, the entire sequence of past tokens must still be stored in memory and used to compute attention scores during generation. Since attention performs a similarity matching of every token representation with every other, it incurs quadratic computational complexity in terms of flops.
@@
There have been recent advances in training LLMs to handle extremely long contexts (up to 1M tokens)Chen et al. (2023); kai (2023); Peng et al. (2023).Such models attain qualitatively new capabilities such as extremely large-scale in-context learning of entire small datasets held in the prompt Reid et al. (2024); Lee et al. (2024); Bertsch et al. (2024). They can also avoid putting multi-modal continuous data through a lossy tokenization scheme Reid et al. (2024); Team (2024) by directly operating at the byte level Xue et al. (2022); Wu et al. (2024). The issue however is that performing inference on such long contexts is very expensive.
@@
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu and Dao (2023); Dao and Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024). Other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory-efficient attention Rabe and Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works, Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. This is the regime of primary interest of this paper. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
@@
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices, instead of linearly in alternatives such as Ring Attention Liu et al. (2023). Our topology aware approach illustrated in Fig. 1. significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
@@
2.  Related Work
Following the ubiquitous success of the transformer architecture, there has been significant effort to mathematically understand the nature and meaning of the attention operation and link it to related architectures, such as recurrent networks and Hopfield Networks. A number of recent works have attempted to study self-attention mathematically through the lens of energy functions. Ramsauer et al. (2020) pioneered this field by performing a similar but distinct analysis to relate self-attention with the modern Hopfield networks Krotov and Hopfield (2016), providing a novel and insightful interpretation of self-attention as performing hetero-associative memory lookups using a high-powered nonlinear similarity function Krotov (2021); Millidge et al. (2022). This work was later extended by Hoover et al. (2023), who derived a modified version of the transformer based off an energy function. Beyond this, other recent work attempted a Bayesian reformulation of attention by deriving a probabilistic generative model which matches the operations performed in a self-attention operation Singh and Buckley (2023).
However, while it has long been known that the softmax operation can be derived as the gradient of the following scalar function:
@@ 2.  Related Work
known as the log-sum-exp, an equivalent function for the self-attention block has not yet been derived.
In particular, consider the energy function proposed in Hoover et al. (2023), which is the logsumexp. Since the gradient in the update rule of that paper is taken with respect to the input to the block, the resulting function is a modified version of the self-attention operation. Similarly, the update rule in Ramsauer et al. (2020) requires the tying of certain weights (K and V) in the attention operation. This restricts the Hopfield derivation to modelling auto-associative retrieval networks, while transformer attention is hetero-associative.
To advance beyond this, our energy function depends on an auxiliary source vector $\zeta$ (see section 4.). Taking the gradient with respect to $\zeta$ yields the exact self-attention operation, as we show below. Consequently, this formulation enables us to express self attention with no weight tying.
@@ 3.  Self-Attention
The self-attention operation can be represented as a set of dot product similarity searches between queries and keys. These similarity scores are then reduced along the sequence axis and softmaxed, so that for a given query, there is a probability distribution of the similarities of each given key. We then take the expectation of the value vectors against this distribution. We denote the queries assigned to a sequence of length $N$ as $\{q_{a},a=1,\cdots N\}$, where each query is a vector of size $d$ that stands for hidden dimension: $q_{a}\in\mathbb{R}^{d}$ and similarly the keys and values $\{(k_{a},v_{a}),a=1,\cdots N\}$. Attention can be written as
@@
The dot product, more explicitly, reads:
@@
The capital Latin indices ($A$) span the hidden dimension $d_{h}$. Naively computing attention in this way requires materializing the $qk$ matrix with computational and memory cost quadratic in the sequence length.
Memory-efficient attention Rabe and Staats (2021) is an iterative way to compute the softmax similarities without ever having to materialize the full attention matrix. It performs the following operations, one query (or a chunk of queries) at a time:
@@
Then, once the values $v$ and softmax denominator $d$ are computed, we divide to get the final softmaxed scores $z^{(j)}=\frac{n^{(j)}}{d^{(j)}}$ for every query index $j$. Computing attention in this iterative manner significantly reduces the required memory.
Flash attention Dao et al. (2022) utilizes a similar approach to reduce the memory and computational cost of attention, but the algorithm is not adapted for multi-GPU computation. Flash attention performs the iterative algorithm of Rabe and Staats (2021) in a blockwise manner, utilizing the block-parallel computational primitives available inside single GPU tensor cores. Additionally, it precisely sizes the blocks such that they can fit into the SRAM of the GPU for the entire attention computation, effectively performing kernel fusion and preventing many unnecessary IO operations.
4.  Self-Attention as a Gradient
Here, we show how the self-attention operation can be written as the gradient of a certain scalar function. In particular, we define a scalar function that depends on the keys, queries, values and additionally, on an auxiliary vector that we refer to as the source $\zeta$. The source is the parameter with respect to which we compute the gradient of the scalar function to obtain the self-attention operation. We need the source in order to write down the generating function of the moments of the distribution above. The source is also the variable with respect to which we can Taylor-expand the generating function and extract the moments as the coefficients of the monomials of $\zeta$ appearing in the Taylor series. Explicitly, we want to find a function $F(q,k,v,\zeta)$ such that:
@@ 4.  Self-Attention as a Gradient
This terminology is inspired by work on energy-based models in machine learning Beal (2003); LeCun et al. (2006); Song and Kingma (2021). A summary of variables and indices is provided in appendix  12.
@@ 4.1.  The energy function
Here, we show how the energy function is given by the cumunlant-generating function associated to the distribution given by attention scores. Taking inspiration from statistical mechanics, where an analogous cumulant-generating function defines the Helmholtz Free energy Wikipedia (2024), we dub our cumulant-generating function the energy function for self-attention.
Let us focus on the case with a single query. As noted above, we leverage the fact that the attention operation can be seen as the computation of the expectation value of the vectors $v$ in the distribution set by the attention scores $z$:
@@
The probability density is given by:
@@
Typically, the denominator or normalization factor is identified with the so-called partition function:
@@
Now, we can compute any moment of the distribution as the $n$-th Taylor coefficient of $Z(\zeta)$ $\forall A_{1},A_{2},\cdots\in\{1,\cdots,d_{h}\}$ :
@@
Therefore, the first moment can be written as:
@@
This quantity is the generating function, a.k.a. the free energy:
@@
To compute causal self-attention, we introduce $N$ sources $\zeta^{i}$ each $\in\mathbb{R}^{d}$ and take
@@
Now, in order to compute the $i$-th element of causal self-attention, we differentiate with respect to $\zeta_{i}$ and set it to zero:
@@
The generalization to the multi-head attention case is straightforward. In this case, there is one key, query and value per head. For $n_{h}$ total heads, the generating function takes the form:
@@
where $W_{O}\in\mathbb{R}^{d_{h}}\times\mathbb{R}^{d_{emb}}$ denotes a head size slice of the output projection weight and $\bar{B}\in\{1,\cdots,d_{h}\}$ spans the intra-head indices. In the index notation above, the head indices are barred whereas the embedding space indices are unbarred. We proceed focusing on the single head case, as it makes the presentation simpler, and the multi-head generalization is immediate. Note that we demonstrate that our energy function approach also can account for safe softmax in Appendix 11.
@@ 4.2.  Bayesian interpretation
The fact that it is possible to derive the self-attention operation as the minimization of an energy function implies that it is possible to provide a Bayesian gloss on self-attention by identifying a likelihood function and showing that we can obtain the forward pass of the attention block from computing the maximum a posteriori estimate of this likelihood.
@@
We denote by $x$ the input to the self-attention block from which we obtain $q,k,v$ from multiplying it by the weights $W_{Q},W_{K},W_{V}$ respectively. Let us minimize the above with respect to $\zeta$ and $z$ simultaneously:
@@ 5.  Tree Attention
In this section we show how the formulation of the attention operation as the gradient of an energy function suggests an efficient parallel strategy for computing it. The key insight is to leverage an efficient algorithm to compute the energy, and then differentiate it in order to obtain an efficient algorithm to compute attention.
@@ 5.1.  Computing the energy efficiently
A crucial fact is that both $\textrm{logsumexp}_{a}$ and $\max_{a}$ are associative operations:
@@ 5.2.  Efficient parallel decoding
This algorithm requires three Allreduce operations in total, meaning that the required time complexity is $O(3(N/p+\log p))$.
@@ 5.3.  Implementation and topology-awareness
(a) Execution time of 16-head Tree Attention vs Ring Attention for different sizes of GPU cluster (from 1 to 16 H100 DGX nodes). Relative execution times are indexed to the Ring Attention times at a sequence length of 80k tokens.
@@
In our empirical experiments , we use Flash Attention 2 Dao (2023) within each device, both for our algorithm and for Ring Attention [^1] A JAX-based Ring Attention implementation that uses Flash Attention 2 can be found here: https://github.com/nshepperd/flash_attn_jax.. We provide a simple JAX implementation of our method in Appendix 9.. Note that our method mirrors Flash Decoding fla (2024) except in that case, the parallelization happens at the level of different streaming multiprocessors (SMs) within a GPU whereas we parallelize between different GPUs. All computations are performed in BF16.
@@ 6.  Results
Similar to Ring Attention, Tree Attention is an exact computation of Attention. Since training and evaluation metrics are the same as for Attention, our experimental results are focused primarily on latency in section 6.1., peak memory usage in section 6.2. and communication volumes in section 6.3.. Since our algorithm computes numerically identical results as the forward pass of standard Attention, our performance results transfer seamlessly to Transformer architectures.
@@ 7.  Discussion and Conclusion
Theoretical asymptotic improvements should be demonstrable in the single-device case, but in practice, this advantage is not realized. SMs in H100 and older architectures communicate through shared global memory, not in a peer-to-peer fashion like individual GPUs in a multi-device setup, making the reduced communication negligible. However, recent experimental instructions in the H100 have enabled peer-to-peer SM communication  NVIDIA (2024), suggesting that these instructions could lead to speedups over Flash Attention on a single device.